# Multi-Agent Reinforcement Learning Simulation for Environmental Policy Synthesis

### v2

@@ 4. Limitations
Primarily, the MARL solution can only be as valid as the underlying simulator; improved optimization solutions cannot transcend the constraints of an imperfect model. This becomes particularly problematic in these environmental domains characterized by unprecedented scenarios or complex emergent phenomena, such as extreme weather patterns as we pass climate tipping points (Farahbakhsh et al., 2024). Current climate simulators may struggle to accurately model the feedback loops between anthropogenic caused rising temperatures and the downstream effects, leading to potentially dangerous blind spots in MARL-based policy recommendations. While there has been progress in open-ended learning within RL frameworks (Wang et al., 2020), this capability must be matched by equally adaptable simulation environments - a requirement that may exceed current modeling capabilities. However, these limitations do not entirely diminish the value of our approach; since our framework is agnostic to the underlying world system model (ESM or IAM or more), it can readily incorporate improved simulators as they become available, ensuring adaptability across various domains and modeling paradigms.

### v3

@@ 4. Limitations
Primarily, the MARL solution can only be as valid as the underlying simulator; improved optimization solutions cannot transcend the constraints of an imperfect model. This becomes particularly problematic in these environmental domains characterized by unprecedented scenarios or complex emergent phenomena, such as extreme weather patterns as we pass climate tipping points (farahbakhsh2024tipping). Current climate simulators may struggle to accurately model the feedback loops between anthropogenic caused rising temperatures and the downstream effects, leading to potentially dangerous blind spots in MARL-based policy recommendations. While there has been progress in open-ended learning within RL frameworks (wang2020enhanced), this capability must be matched by equally adaptable simulation environments - a requirement that may exceed current modeling capabilities. However, these limitations do not entirely diminish the value of our approach; since our framework is agnostic to the underlying world system model (ESM or IAM or more), it can readily incorporate improved simulators as they become available, ensuring adaptability across various domains and modeling paradigms.
@@ 5. Conclusion
{acks}
James Rudd-Jones is supported by grants from the UK EPSRC-DTP (Award 2868483). Mirco Musolesi was supported by the Italian Ministry of University and Research (MUR) through the project PRIN 2022 “Machine-learning based control of complex multi-agent systems for search and rescue operations in natural disasters (MENTOR)” funded by the European Union - NextGenerationEU.